# Creation of Negatively Charged GeV and SnV centers in Nanodiamonds via Ion Implantation

### v1

@@ Abstract
tin vacancy centers, germanium vacancy centers, nanodiamonds, ion implantation, single photonSelene Sachero1,
Richard Waltrich1,
Emilio Corte2,
Sviatoslav Ditalia Tchernij2,
Alexander Kubanek1*
1 Institute for Quantum Optics, Ulm University, 89081 Ulm, Germany
@@
*alexander.kubanek@uni-ulm.de
@@ 3.1 Spectroscopy of GeV- centers
The histogram in figure 3a) illustrates the inhomogeneous distributions of the ZPL for all measured GeV- centers. The data are distributed according to a normal distribution centered at nm. In figure 3b), we show that, despite the high implantation fluence ( ), we were able to observe the fine structure of an isolated GeV- center. C and D transitions were fitted with a Lorentzian curve and the resulting wavelengths are nm and nm, respectively. This results in a GS splitting of 428 GHz which indicates large strain in the surrounding ND lattice which originates from the damages induced by the high implantation fluence. We proceed with the optical characterization via resonant excitation. We selected a GeV- center from the cluster in figure 3c). This emitter exhibited a resonant wavelength of (602.16 0.02) and a linewidth of (440 40 ) MHz (figure 3d)), which is four times broader than the lifetime limited linewidth reported for bulk diamond [34]. We address the remaining broadening to spectral diffusion which was observable during the PLE scans, and it is consistent with the persistent blinking fluorescence under CW resonant drive (see section 3 in the SI). To prove the quantum nature of the measured GeV- center CW resonant autocorrelation measurement was performed. The antibunching behavior results in a minimum value of . No background correction was performed on the data, so the high purity was a result of the resonant drive.
@@ 3.2 Spectroscopy of SnV- centers
The histogram in figure 4a) illustrates the inhomogeneous distributions of the ZPL for all measured SnV- centers. The data are distributed according to a Gaussian distribution centered at nm. Figure 4b) presents an example of the PL spectrum of an isolated emitter. The fine structure of the emitter is evident; C and D transition lines were fitted with a Lorentzian and the resulted wavelengths are nm and nm, respectively. This results in a GS splitting of GHz. The latter value is in accordance with the predicted theoretical value of 850 MHz [62] which indicates that the surrounding ND lattice has low strain. We proceed with the optical characterization via resonant excitation. We selected a single SnV- center from the cluster shown in figure 4c). This emitter exhibited a resonant wavelength of nm and a linewidth of MHz as shown in figure 4d). The measured linewidth was found to be significantly broadened compared to the lifetime-limited linewidth of 30 MHz reported for bulk diamond [63]. The broadening is attributed to spectral diffusion, which was observed during the PLE scans. The broadening is consistent with the persistent blinking fluorescence observed under CW resonant drive (see section 2.2 in the SI). Despite the instability under resonant excitation, the emitter exhibited considerable brightness. Spectral diffusion can be attributed to a charge instability, which can be mitigated by oxygen annealing at 450 or by adding a CW 445 nm laser used to stabilize the charge [64]. The strong antibunching behavior (figure 4e)), proved the high purity of the single-photon emission under resonant excitation. No background correction was performed on the data so the high purity was a result of the resonant drive.

### v2

@@ Abstract
tin vacancy centers, germanium vacancy centers, nanodiamonds, ion implantation, single photonSelene Sachero1, Richard Waltrich1, Emilio Corte2, Jens Fuhrmann1, Sviatoslav Ditalia Tchernij2, Fedor Jelezko1,3, Alexander Kubanek1,3*
1 Institute for Quantum Optics, Ulm University, Albert-Einstein-Allee 11, 89081 Ulm, Germany
@@
3 Center for Integrated Quantum Science and Technology (IQst), Ulm University, Albert-Einstein-Allee 11, 89081, Ulm, Germany
*Corresponding author: alexander.kubanek@uni-ulm.de
@@ 3.1 Spectroscopy of GeV- centers
The histogram in figure 3a) illustrates the inhomogeneous distributions of the ZPL for all measured GeV- centers. The data are distributed according to a normal distribution centered at nm. In figure 3b), we show that, despite the high implantation fluence ( ), we were able to observe the fine structure of an isolated GeV- center. C and D transitions were fitted with a Lorentzian curve and the resulting wavelengths are nm and nm, respectively. This results in a GS splitting of (430 20) GHz which indicates large strain in the surrounding ND lattice which originates from the damages induced by the high implantation fluence. We proceed with the optical characterization via resonant excitation. We selected a GeV- center from the cluster in figure 3c). This emitter exhibited a resonant wavelength of (602.16 0.02) and a linewidth of (440 40 ) MHz (figure 3d)), which is four times broader than the lifetime limited linewidth reported for bulk diamond [34]. We address the remaining broadening to spectral diffusion which was observable during the PLE scans, and it is consistent with the persistent blinking fluorescence under CW resonant drive (see section 3 in the SI). To prove the quantum nature of the measured GeV- center CW resonant autocorrelation measurement was performed. The antibunching behavior results in a minimum value of . No background correction was performed on the data, so the high purity was a result of the resonant drive.
@@ 3.2 Spectroscopy of SnV- centers
The histogram in figure 4a) illustrates the inhomogeneous distributions of the ZPL for all measured SnV- centers. The data are distributed according to a Gaussian distribution centered at nm. Figure 4b) presents an example of the PL spectrum of an isolated emitter. The fine structure of the emitter is evident; C and D transition lines were fitted with a Lorentzian and the resulted wavelengths are nm and nm, respectively. This results in a GS splitting of GHz. The latter value is in accordance with the predicted theoretical value of 850 GHz [62] which indicates that the surrounding ND lattice has low strain. We proceed with the optical characterization via resonant excitation. We selected a single SnV- center from the cluster shown in figure 4c). This emitter exhibited a resonant wavelength of nm and a linewidth of MHz as shown in figure 4d). The measured linewidth was found to be significantly broadened compared to the lifetime-limited linewidth of 30 MHz reported for bulk diamond [63]. The broadening is attributed to spectral diffusion, which was observed during the PLE scans. The broadening is consistent with the persistent blinking fluorescence observed under CW resonant drive (see section 2.2 in the SI). Despite the instability under resonant excitation, the emitter exhibited considerable brightness. Spectral diffusion can be attributed to a charge instability, which can be mitigated by oxygen annealing at 450 or by adding a CW 445 nm laser used to stabilize the charge [64]. The strong antibunching behavior (figure 4e)), proved the high purity of the single-photon emission under resonant excitation. No background correction was performed on the data so the high purity was a result of the resonant drive.